# Revealing the nature of $\Omega_{c}$-like states from pentaquark perspective

### v2

@@ III Numerical results and discussion
Electromagnetic form factors, which describe the response of composite particles to electromagnetic surveys, provide an important tool for understanding the structure of bound states in QCD. The magnetic dipole moment is the leading-order response of a bound system to a weak external magnetic field. It therefore offers an exquisite laboratory for studying the internal structure of composite particles, which is governed by the quark-gluon dynamics of QCD. Therefore, we systematically study the magnetic dipole moment of controversial states whose internal structure is not elucidated and we try to offer a different point of view to unravel the internal structure of these states. Inspired by the $\Omega_{c}$ states observed by the LHCb Collaboration, we focus on the scenario of the diquark-diquark-antiquark pentaquark interpretation of the $\Omega_{c}$-like states with both $J^{P}=\frac{1}{2}^{-}$ and $J^{P}=\frac{3}{2}^{-}$ quantum numbers, and study the electromagnetic properties of these states with the QCD light-cone sum rules by employing the distribution amplitudes of the photon. From the obtained numerical results, we conclude that the magnetic dipole moments of the $\Omega_{c}$-like states can reflect their inner structures, which can be used to distinguish their spin-parity quantum numbers. The order of the magnetic dipole moments obtained for these states implies that they are accessible by experiments. Measuring the magnetic dipole moments of the $\Omega_{c}$-like states in future experimental facilities can be very helpful for understanding the internal organization and identifying the quantum numbers of these states. We hope that in the future we will have more experimental data on a variety of physical quantities of these states.
Appendix: Magnetic dipole moments of the spin-1/2 $\Omega_{c}$ states
$\mu_{\Omega_{c}}$ $=\frac{e^{\frac{m^{2}_{\Omega_{c}}}{\rm{M^{2}}}}}{\,\lambda^{2}_{\Omega_{c}}}\Bigg{\{}-\frac{e_{q}}{2^{13}\times 3\times 5m_{c}^{2}\pi^{8}}\Bigg{[}2m_{c}^{14}I[-7,5]+5m_{q_{3}}m_{c}^{13}I[-6,4]+m_{c}^{12}\big{(}-10m_{q_{1}}m_{q_{2}}I[-6,4]+8I[-6,5]\big{)}$
@@
$+\frac{P_{1}P_{5}}{2^{15}3^{4}m_{c}^{2}\pi^{6}}\Bigg{[}-24e_{q_{3}}m_{c}^{3}I_{3}[\mathcal{\tilde{S}}]I[-1,1]-(e_{q_{1}}+2e_{q_{12}}+e_{q_{2}})f_{3\gamma}\pi^{2}I_{2}[\mathcal{V}]\Big{(}41m_{c}^{3}\ I[-1,0]+26m_{q_{3}}(m_{c}^{4}I[-2,0]-I[0,0])\Big{)}$
@@
$+52m_{c}^{2}I[-1,1]+208I[0,1]\big{)}I_{3}[\mathcal{S}]\Big{)}-32(e_{q_{1}}+2e_{q_{12}}+e_{q_{2}})f_{3\gamma}\pi^{2}\Big{(}2m_{c}^{3}I[-1,0]+m_{q_{3}}(m_{c}^{4}I[-2,0]-I[0,0])-m_{c}I[0,0]$
$+m_{c}^{5}(I[-2,0]-2I[2,0])\Big{)}\psi^{a}[u_{0}]+16\chi e_{q_{3}}m_{c}^{3}\Big{(}2m_{c}^{6}I[-4,2]-3m_{c}^{4}I[-3,2]+I[-1,2]\Big{)}\varphi_{\gamma}[u_{0}]\Bigg{]}$
@@
where $\mathcal{F}$ represents the corresponding photon DAs.
@@
     Parameters      $ssuc\bar{u}$      $susc\bar{u}$      $ussc\bar{u}$      $ssdc\bar{d}$      $sdsc\bar{d}$      $dssc\bar{d}$
     $e_{q_{1}}$      $e_{s}$      $e_{s}$      $e_{u}$      $e_{s}$      $e_{s}$      $e_{d}$
     $e_{q_{2}}$      $e_{s}$      $e_{u}$      $e_{s}$      $e_{s}$      $e_{d}$      $e_{s}$
     $e_{q_{3}}$      $e_{u}$      $e_{s}$      $e_{s}$      $e_{d}$      $e_{s}$      $e_{s}$
     $e_{q_{12}}$      $e_{s}$      --      --      $e_{s}$      --      --
     $e_{q_{13}}$      --      $e_{s}$      --      --      $e_{s}$      --
     $e_{q_{23}}$      --      --      $e_{s}$      --      --      $e_{s}$
     $m_{q_{1}}$      $m_{s}$      $m_{s}$      $m_{u}$      $m_{s}$      $m_{s}$      $m_{d}$
     $m_{q_{2}}$      $m_{s}$      $m_{u}$      $m_{s}$      $m_{s}$      $m_{d}$      $m_{s}$
     $m_{q_{3}}$      $m_{u}$      $m_{s}$      $m_{s}$      $m_{d}$      $m_{s}$      $m_{s}$
     $m_{q_{12}}$      $m_{s}$      --      --      $m_{s}$      --      --
     $m_{q_{13}}$      --      $m_{s}$      --      --      $m_{s}$      --
     $m_{q_{23}}$      --      --      $m_{s}$      --      --      $m_{s}$
     $P_{3}$      $\langle\bar{s}s\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{q}q\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{q}q\rangle$
     $P_{4}$      $\langle\bar{s}s\rangle$      $\langle\bar{q}q\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{q}q\rangle$      $\langle\bar{s}s\rangle$
     $P_{5}$      $\langle\bar{q}q\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{q}q\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{s}s\rangle$
     $P_{6}$      $\langle\bar{s}s\rangle$      --      --      $\langle\bar{s}s\rangle$      --      --
     $P_{7}$      --      $\langle\bar{s}s\rangle$      --      --      $\langle\bar{s}s\rangle$      --
     $P_{8}$      --      --      $\langle\bar{s}s\rangle$      --      --      $\langle\bar{s}s\rangle$

### v3

@@ III Numerical results
Electromagnetic form factors, which describe the response of composite particles to electromagnetic surveys, provide an important tool for understanding the structure of bound states in QCD. The magnetic dipole moment is the leading-order response of a bound system to a weak external magnetic field. It therefore offers an exquisite laboratory for studying the internal structure of composite particles, which is governed by the quark-gluon dynamics of QCD. Therefore, we systematically study the magnetic dipole moment of controversial states whose internal structure is not elucidated and we try to offer a different point of view to unravel the internal structure of these states. Inspired by the $\Omega_{c}$ states observed by the LHCb Collaboration, we focus on the scenario of the diquark-diquark-antiquark pentaquark interpretation of the $\Omega_{c}$ states with both $J^{P}=\frac{1}{2}^{-}$ and $J^{P}=\frac{3}{2}^{-}$ quantum numbers, and study the electromagnetic properties of these states with the QCD light-cone sum rules by employing the distribution amplitudes of the photon. From the obtained numerical results, we conclude that the magnetic dipole moments of the $\Omega_{c}$ states can reflect their inner structures, which can be used to distinguish their spin-parity quantum numbers. The order of the magnetic dipole moments obtained for these states implies that they are accessible by experiments. Measuring the magnetic dipole moments of the $\Omega_{c}$ states in future experimental facilities can be very helpful for understanding the internal organization and identifying the quantum numbers of these states. We hope that in the future we will have more experimental data on a variety of physical quantities of these states.
Figure 1: Dependencies of the magnetic dipole moments of $\Omega_{c}$ states on $\rm{M^{2}}$ at three values of $\rm{s_{0}}$; (a), (c), and (e) for spin-$\frac{1}{2}$ states, and; (b), (d), and (f) for spin-$\frac{3}{2}$ states.
Appendix: Analytical results of the magnetic dipole moments of the spin-1/2 $\Omega_{c}$ states
$\mu_{\Omega_{c}}$ $=\frac{e^{\frac{m^{2}_{\Omega_{c}}}{\rm{M^{2}}}}}{\,\lambda^{2}_{\Omega_{c}}}\Bigg{\{}-\frac{e_{d}}{2^{13}\times 3\times 5m_{c}^{2}\pi^{8}}\Bigg{[}2m_{c}^{14}I[-7,5]+5m_{q_{3}}m_{c}^{13}I[-6,4]+m_{c}^{12}\big{(}-10m_{q_{1}}m_{q_{2}}I[-6,4]+8I[-6,5]\big{)}$
@@
$+\frac{P_{1}P_{5}}{2^{15}3^{4}m_{c}^{2}\pi^{6}}\Bigg{[}-24e_{q_{3}}m_{c}^{3}I_{3}[\mathcal{\tilde{S}}]I[-1,1]-(e_{q_{1}}+2e_{q_{12}}+e_{q_{2}})f_{3\gamma}\pi^{2}I_{2}[\mathcal{V}]\Big{(}41m_{c}^{3}\ I[-1,0]+26m_{q_{3}}(m_{c}^{4}I[-2,0])\Big{)}$
@@
$+52m_{c}^{2}I[-1,1]+208I[0,1]\big{)}I_{3}[\mathcal{S}]\Big{)}-32(e_{q_{1}}+2e_{q_{12}}+e_{q_{2}})f_{3\gamma}\pi^{2}\Big{(}2m_{c}^{3}I[-1,0]+m_{q_{3}}(m_{c}^{4}I[-2,0]-I[0,0])$
$-m_{c}I[0,0]+m_{c}^{5}(I[-2,0]-2I[2,0])\Big{)}\psi^{a}[u_{0}]+16\chi e_{q_{3}}m_{c}^{3}\Big{(}2m_{c}^{6}I[-4,2]-3m_{c}^{4}I[-3,2]+I[-1,2]\Big{)}\varphi_{\gamma}[u_{0}]\Bigg{]}$
@@
where $\mathcal{M}=m_{c}+2m_{s}$, and $\mathcal{F}$ represents the corresponding photon DAs.
@@
     Parameters      $ssdc\bar{d}$      $sdsc\bar{d}$      $dssc\bar{d}$
     $e_{q_{1}}$      $e_{s}$      $e_{s}$      $e_{d}$
     $e_{q_{2}}$      $e_{s}$      $e_{d}$      $e_{s}$
     $e_{q_{3}}$      $e_{d}$      $e_{s}$      $e_{s}$
     $e_{q_{12}}$      $e_{s}$      --      --
     $e_{q_{13}}$      --      $e_{s}$      --
     $e_{q_{23}}$      --      --      $e_{s}$
     $m_{q_{1}}$      $m_{s}$      $m_{s}$      $m_{d}$
     $m_{q_{2}}$      $m_{s}$      $m_{d}$      $m_{s}$
     $m_{q_{3}}$      $m_{d}$      $m_{s}$      $m_{s}$
     $m_{q_{12}}$      $m_{s}$      --      --
     $m_{q_{13}}$      --      $m_{s}$      --
     $m_{q_{23}}$      --      --      $m_{s}$
     $P_{3}$      $\langle\bar{s}s\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{q}q\rangle$
     $P_{4}$      $\langle\bar{s}s\rangle$      $\langle\bar{q}q\rangle$      $\langle\bar{s}s\rangle$
     $P_{5}$      $\langle\bar{q}q\rangle$      $\langle\bar{s}s\rangle$      $\langle\bar{s}s\rangle$
     $P_{6}$      $\langle\bar{s}s\rangle$      --      --
     $P_{7}$      --      --      $\langle\bar{s}s\rangle$
     $P_{8}$      --      --      $\langle\bar{s}s\rangle$